The cost of living in a certain city has been increasing so much that an item that cost $1 today would cost (1.08)^t dollars in t years. Using this relationship, how much did today's one-dollar item cost 9 years ago?
Answer by stanbon(75887)

The cost of living in a certain city has been increasing so much that an item that cost $1 today would cost (1.08)^t dollars in t years. Using this relationship, how much did today's one-dollar item cost 9 years ago?
----------
Let "x" be the value 9 years ago.
EQUATION:
x(1.08)^9=$1
x= 1/(1.08)^9
x = 1/1.999004627...
x= $0.50 or 50 cents
--------
Cheers,
Stan H.